OpenStudy (anonymous):

The equation for the chemical reaction shown is not balanced. What number should replace the question mark to balance this equation? 2Al + ?Cl2 → 2AlCl3

OpenStudy (anonymous):

A. the number 3 B. the number 2 C. the number 5 D. the number 4

OpenStudy (anonymous):

@thomaster

thomaster (thomaster):

On the left side of the arrow, we have 2 aluminium atoms, same for the right side of the arrow. how many chlorine atoms do we have on the left site of the arrow? and how about the right side? You need to put the number on the ? that will make the number of chlorine atoms equal on both sides of the arrow.

OpenStudy (anonymous):

3?

thomaster (thomaster):

that's correct :) the number behind an atom means that there are 2 of the atom the number before the molecule will multiply every atom in the molecule by that number. So if you have \(\sf\large H_2O\), it means we have 2 hydrogen atoms and 1 oxygen atom. \(\sf\large 2H_2O\) means we have 4 hydrogen atoms (2x2) and 2 oxygen atoms.

OpenStudy (anonymous):

But than won't 3*2=6CI?

thomaster (thomaster):

yes, and on the right side of the arrow we have 2AlCl3. the 2 in front of that molecule means that Cl3 will be multiplied by 2. And 3*2=6 Cl atoms

OpenStudy (anonymous):

But 2 is also on the other side so won't 6*2= 12 making it uneven?

thomaster (thomaster):

\(\Large\sf 2Al ~~+ ~~?Cl_2 → 2AlCl_3\) \(\Large\sf 2AlCl_3=2\ aluminium\ atoms\ and\ 2*3=6\ chlorine\ atoms \) \(\Large\sf ?Cl_2 =x\ times\ 2\ chlorine\ atoms\) when you put 3 instead of x, you'd have 3 times 2 = 6 chlorine atoms on the right side we also have 6 chlorine atoms, so they're equal now

OpenStudy (anonymous):

Ohh!! Thanks

OpenStudy (unklerhaukus):